# Black-box Adversaries from Latent Space: Unnoticeable Attacks on Human Pose and Shape Estimation

### v1

@@ Abstract
Expressive human pose and shape (EHPS) estimation is vital for digital human generation, particularly in live-streaming applications. However, most existing EHPS models focus primarily on minimizing estimation errors, with limited attention on potential security vulnerabilities. Current adversarial attacks on EHPS models often require white-box access (e.g., model details or gradients) or generate visually conspicuous perturbations, limiting their practicality and ability to expose real-world security threats. To address these limitations, we propose a novel Unnoticeable Black-Box Attack (UBA) against EHPS models. UBA leverages the latent-space representations of natural images to generate an optimal adversarial noise pattern and iteratively refine its attack potency along an optimized direction in digital space. Crucially, this process relies solely on querying the model’s output, requiring no internal knowledge of the EHPS architecture, while guiding the noise optimization toward greater stealth and effectiveness. Extensive experiments and visual analyses demonstrate the superiority of UBA. Notably, UBA increases the pose estimation errors of EHPS models by 17.27%–58.21% on average, revealing critical vulnerabilities. These findings underscore the urgent need to address and mitigate security risks associated with digital human generation systems.
1 Introduction
@@
Expressive human pose and shape (EHPS) estimation from monocular images or videos is fundamental to digital human modeling and underpins key applications such as live streaming, virtual reality (VR), and interactive gaming [26, 31, 46, 42, 2, 13, 12]. In these scenarios, EHPS systems play a critical role in animating realistic digital avatars capable of dynamically replicating human movements and facial expressions in real time. Although recent advancements have significantly enhanced estimation accuracy, the security vulnerabilities of such models remain largely underexplored. Growing evidence [39, 3, 33, 45] indicate that these systems are susceptible to adversarial perturbations; however, existing attack methods against EHPS [23] often depend on access to internal model parameters or gradient information and typically produce visually noticeable artifacts. These limitations not only hinder the practical deployment of such attacks but also contribute to the limited attention this issue has received within the research community.
@@
Existing EHPS studies primarily focus on reducing errors in body pose estimation [46, 42] and facial pose estimation [18, 28, 24, 2]. This is typically achieved using parametric human body models such as SMPL-X [31] or SMPLer-X [2], which leverage deep neural networks to capture the complex geometry and motion of the human body. While these models exhibit strong performance, their robustness against adversarial attacks has received limited attention. TBA [23] was the first to propose a white-box transferable adversarial attack against EHPS. It utilizes the gradient of the SMPLer-X model to generate adversarial examples that significantly increase the estimation error across various EHPS models, thereby exposing potential security vulnerabilities. However, the perturbations produced by TBA are visually perceptible and fail to pass human inspection, which limits their applicability in real-world scenarios. Despite their effectiveness, such attacks have yet to attract sufficient attention from the research community. Considering these factors, we argue that an ideal adversarial attack on EHPS should meet three key criteria: i) the perturbation must be visually imperceptible; ii) the required prior knowledge of the attacker should be minimal, aligning with realistic threat models; iii) the adversarial examples must achieve a high error growth rate. However, a substantial gap remains between current methods and these ideal standards.
@@
To address this gap, we propose a novel Unnoticeable Black-box Attack (UBA) targeting EHPS models. Inspired by [6, 32], UBA perturbs the latent space of a pretrained variational autoencoder (VAE) [20, 11] by injecting low-magnitude Gaussian noise. Leveraging the nonlinear generative capabilities of the decoder of VAE [19, 37], this will produce an initial adversarial example that is visually indistinguishable from the original input. The resulting pixel-wise difference is treated as an initial perturbation and subsequently refined through an iterative optimization process guided by a customized multi-task loss function. This loss simultaneously maximizes the discrepancy in the EHPS model’s predictions while minimizing perceptual deviation from the clean image. Notably, the attack is conducted under a strict query budget, ensuring both query efficiency and imperceptibility. By exploiting latent-space amplification and aligning perturbations with task-relevant features, UBA achieves high attack efficacy while maintaining visual stealth. In conclusion, our main contributions are:
$\bullet$ We propose the first black-box adversarial attack tailored to EHPS systems, satisfying realistic constraints and highlighting critical vulnerabilities for digital human generation.
$\bullet$ By injecting noise into the VAE latent space, we generate visually imperceptible adversarial examples that exploit the decoder’s nonlinear generative capabilities.
$\bullet$ A custom multi-task loss balances attack strength and imperceptibility, with iterative refinement performed under a strict query limit to avoid detection.
$\bullet$ Extensive experiments demonstrate that our method significantly degrades pose estimation performance, increasing EHPS models’ pose estimation errors by 17.27%–58.21% on average.
2 Related Work
@@
Expressive Human Pose and Shape Estimation. Expressive Human Pose and Shape Estimation (EHPS) is fundamental to digital human modeling, with applications in virtual environments, live broadcasting, avatars, and animation [40, 22, 30, 47, 29, 34]. Recent years have seen significant advances in improving the accuracy and realism of EHPS models. Pavlakos et al. [31] introduced SMPL-X, a comprehensive 3D model incorporating articulated hands and expressive facial features for improved monocular pose capture. Moon et al. [28] proposed Hand4Whole, leveraging MCP joint features to predict wrist rotations precisely while excluding full-body context for better finger estimation. Lin et al. [24] developed OSX, a one-stage framework using a Component Aware Transformer to reconstruct whole-body meshes without post-processing. Cai et al. [2] presented SMPLer-X, a generalist foundation model trained on diverse datasets, achieving strong generalization and competitive performance across EHPS benchmarks. However, existing research has predominantly concentrated on enhancing estimation accuracy, with limited attention to the models’ vulnerability under attack threats.
@@
Adversarial Attack against Deep Neural Networks. Adversarial attacks are commonly employed to assess the robustness of deep neural networks (DNNs) against malicious inputs [16, 9, 34, 14, 15, 44]. These attacks introduce subtle perturbations that mislead models while preserving the visual appearance of the input. Huang et al. [17] proposed TT3D, which integrates NeRF-based multi-view reconstruction with dual optimization of texture features and MLP parameters to generate natural-looking, transferable 3D adversarial examples. Chen et al. [5] presented the first content-aware, unrestricted adversarial attack based on diffusion models, enhancing transferability and realism by optimizing semantic perturbations in latent space. Chen et al. [4] proposed DiffAttack, a diffusion-based framework that achieves highly transferable and imperceptible black-box attacks by jointly modulating cross-attention and self-attention mechanisms. Despite these advancements, the robustness of EHPS models remains largely underexplored. Li et al. [23] proposed TBA consisting of a dual heterogeneous noise generator, the first EHPS-targeted attack; however, its visible perturbations and reliance on gradient access limit its practicality. Consequently, TBA has failed to draw sufficient attention from the research community regarding EHPS security risks. To address this gap, we propose a unnoticeable black-box attack that reveals critical vulnerabilities in EHPS models.
3 Preliminaries
Threat Model. This paper investigates black-box adversarial attack techniques targeting EHPS models. Specifically, the attacker lacks access to any details of the model’s training process, including the dataset, architecture, and parameters, and is only permitted a limited number of queries via the model’s API. The generated adversarial examples are directly submitted to the model for inference. We aim to raise awareness in the research community about security risks in the domain of digital humans. Other attack types with limited practical applicability (e.g., training-controlled or model modification attacks) are excluded from consideration, as they require additional attacker capabilities and are thus unsuitable for this study.
The Main Pipeline of EHPS. Expressive Human Pose and Shape (EHPS) estimation aims to reconstruct detailed 3D representations of the human body, face, and hands from monocular images. This capability underpins a wide array of applications, including virtual environments, live broadcasting, avatar generation, and animation. Formally, the ground-truth pose is denoted as $\alpha\in\Phi^{53\times 3}$, comprising the body pose $\alpha_{\text{body}}\in\Phi^{22\times 3}$, left- and right-hand poses $\alpha_{\text{lhand}},\alpha_{\text{rhand}}\in\Phi^{15\times 3}$, and the jaw pose $\alpha_{\text{jaw}}\in\Phi^{1\times 3}$. Additionally, individual-specific geometry and facial expressions are encoded by the shape vector $\beta\in\Phi^{10}$ and the expression vector $\gamma\in\Phi^{10}$, respectively. Let $\mathcal{P}$ denote the EHPS model; given a unseen monocular image $x^{{}^{\prime}}$, the objective is to estimate the complete set of pose, shape, and expression parameters $\left\{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right\}$, e.g., $\mathcal{P}(x^{{}^{\prime}})=\left\{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right\}$. The estimation process minimizes the parameter reconstruction loss:
$\min_{{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}}}\left(\|\alpha^{{}^{\prime}}-\alpha\|+\|\beta^{{}^{\prime}}-\beta\|+\|\gamma^{{}^{\prime}}-\gamma\|\right).$ (1)
Finally, 3D joints are rendered using the SMPL-X function $\mathcal{S}$ and a joint regressor $\mathcal{R}$, denoted as $\mathcal{R}(\mathcal{S}(\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}))$.
4 Unnoticeable Black-box Attack (UBA)
4.1 Problem Formulation
As noted in previous sections, existing EHPS models primarily aim to reduce human pose and shape estimation errors, with limited emphasis on their vulnerability to attacks. Although prior work [23] introduced the first adversarial attack against EHPS, exposing its inherent fragility, its real-world applicability remains constrained due to the visible nature of adversarial perturbations and the unrealistic assumption that attackers have access to internal model information. These limitations hinder the community’s full recognition of the intrinsic risks associated with EHPS. To address this, we propose an unnoticeable black-box attack tailored to a realistic setting, where the attacker can only interact with the model through query outputs. To avoid detection by the owners of the EHPS model due to excessive queries, the attack is carried out under a strict query budget.
Attacker’s Goals. The UBA is designed with two primary objectives: 1) effectiveness and 2) unnoticeability. Specifically, the effectiveness demands that the adversarial examples significantly increase errors in human pose and shape estimation. The unnoticeability ensures that EHPS model owners are unable to identify the adversarial examples. The formalization is as follows:
$\mathcal{P}(\hat{x})=\left\{\hat{\alpha}\in\Phi^{53\times 3},\hat{\beta}\in\Phi^{10},\hat{\gamma}\in\Phi^{10}\right\},$ (2)
$\max_{{\hat{\alpha},\hat{\beta},\hat{\gamma}}}\left(\|\hat{\alpha}-\alpha\|+\|\hat{\beta}-\beta\|+\|\hat{\gamma}-\gamma\|\right),\quad\text{s.t.}\quad\left\|x-\hat{x}\right\|\leq\varepsilon,$ (3)
where $x$ denotes a clean image, $x^{adv}$ represents the corresponding adversarial sample, and $\varepsilon$ indicates the perturbation threshold. $\left\{\alpha,\beta,\gamma\right\}$ denotes the ground-truth human pose and shape parameters.
@@ 4.1 Problem Formulation
Figure 1: Overview of the Unnoticeable Black-box Attack (UBA) pipeline. The framework consists of two stages: a) noise injection in latent space, and b) noise enhancement.
4.2 The proposed Method
As shown in Fig. 1, our proposed UBA method is a black-box attack framework designed for EHPS models, consisting of two primary stages: 1) injecting random noise into the latent space; and 2) optimizing the noise to maximize attack effectiveness while remaining unnoticeable. Notably, our goal is to reveal vulnerabilities in EHPS models and raise awareness of their security risks, not to cause harm.
@@ 4.2 The proposed Method
Noise Injection in Latent Space. Motivated by the insights of Shukla et al. [35], who challenged the necessity of norm-bounded constraints in pixel space, we argue that adversarial perturbations are more effective when applied in the latent space. Unlike conventional methods that manipulate input images directly [8, 27, 5, 4], this strategy targets internal representations that more directly govern model predictions. Importantly, perturbing the latent space allows for adversarial behavior while preserving high perceptual fidelity in the image space. To reduce attack-side burden and increase real-world applicability, we directly utilize a publicly available pretrained generative model, without any additional training or fine-tuning. To this end, we employ an efficient, pretrained Variational Autoencoder (VAE) [20] to generate adversarial samples.
@@
The VAE is trained by maximizing the evidence lower bound and utilizes the reparameterization trick to enable differentiable sampling:
$\mathcal{L}(\phi,\theta;x)=\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{\mathrm{KL}}(q_{\phi}(z|x),p(z)),$ (4)
@@
where $\phi$ and $\theta$ denote the encoder and decoder parameters, respectively. $p(z)=\mathcal{N}(0,I)$ is the standard Gaussian prior. $\mathcal{G}_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n},z\mapsto x$, $d$ and $n$ denote the dimension of the latent space and the pixel space. This design yields a smooth and continuous latent space [21].
Building on this property, we inject an additional Gaussian noise $\epsilon_{2}$ into the latent representation $z$, creating a variationally grounded sampling-based perturbation that preserves the Gaussianity of the distribution. Specifically, we define $\hat{z}=z+\eta\cdot\epsilon_{2}$, where $\eta$ is a hyperparameter controlling the perturbation magnitude, so that $p(\hat{z})\sim\mathcal{N}(0,I)$. When decoded by the Decoder $\mathcal{G}_{\theta}$, the perturbed latent representation $\hat{z}$ yields outputs $\hat{x}=\mathcal{G}_{\theta}(\hat{z})$. We perform a first-order Taylor expansion of $\mathcal{G}_{\theta}(z)$:
$\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2})=\mathcal{G}_{\theta}(z)+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})+o(||\eta\cdot\epsilon_{2}||),$ (6)
where $J_{\mathcal{G}_{\theta}}(z)$ denotes the Jacobian of the decoder $\mathcal{G}_{\theta}$ at $z$. By omitting the higher-order terms $o(||\eta\cdot\epsilon_{2}||)$, we can obtain:
@@
$\Delta x\approx J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\leq\eta\cdot||J_{\mathcal{G}_{\theta}}(z)||_{2}\cdot||\epsilon_{2}||.$ (9)
@@
$\mathcal{P}(\hat{x})=\mathcal{P}(x+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}))\approx\mathcal{P}(x)+J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (10)
We define the threshold for significant changes in the estimation parameters of the EHPS model as $\delta$. If the deviation exceeds this threshold, the model is considered to be misled, i.e., $||\mathcal{P}(\hat{x})-\mathcal{P}(x)||\geq\delta$. According to Eq. 10, this leads to the inequality $J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\geq\delta$.
By combining the Jacobian matrices into a single term $\mathbb{J}=J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)$, the inequality can be reformulated as $||\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||\geq\delta$. Given that $\epsilon_{2}\sim\mathcal{N}(0,I)$, the expected perturbation is:
$\mathbb{E}[||\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||^{2}_{2}]=\eta^{2}\mathbb{E}[\epsilon_{2}^{\top}\mathbb{J}^{\top}\mathbb{J}\epsilon_{2}]=\eta^{2}\text{tr}(\mathbb{J}^{\top}\mathbb{J}).$ (11)
@@
$||\mathbb{J}\cdot(\eta\cdot\upsilon_{1})||_{2}=\eta\cdot\sigma_{1}\geq\delta\Rightarrow\eta\geq\frac{\delta}{\sigma_{1}}.$ (13)
@@
$\hat{z}^{(t+1)}=\hat{z}^{(t)}+\xi_{z}\nabla_{\hat{z}^{(t)}}\mathcal{L}(\mathcal{P}(\mathcal{G}_{\theta}(\hat{z}^{(t)})),(\alpha,\beta,\gamma)),$ (14)
@@
$\nabla_{\hat{z}^{(t)}}\mathcal{L}(\mathcal{P}(\mathcal{G}_{\theta}(\hat{z}^{(t)})),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial\mathcal{G}_{\theta}}\frac{\partial\mathcal{G}_{\theta}}{\partial\hat{z}^{(t)}},$ (15)

### v2

# LatentStealth: Unnoticeable and Efficient Adversarial Attacks on Expressive Human Pose and Shape Estimation

@@ Abstract
Expressive human pose and shape estimation (EHPS) plays a central role in digital human generation, particularly in live-streaming applications. However, most existing EHPS models focus primarily on minimizing estimation errors, with limited attention on potential security vulnerabilities, such as generating inappropriate content, violent actions, or racially offensive gestures and expressions. Current adversarial attacks on EHPS models often generate visually conspicuous perturbations, limiting their practicality and ability to expose real-world security threats. To address this limitation, we propose an unnoticeable adversarial method, termed LatentStealth, specifically tailored for EHPS models. The key idea is to exploit the structured latent representations of natural images as the medium for crafting perturbations. Instead of injecting noise directly into the pixel space, our method projects inputs into the latent space, where adversarial patterns are generated and progressively refined along optimized directions. This latent-space manipulation enables the attack to maintain high imperceptibility while preserving its effectiveness. Furthermore, as the optimization process is guided by only a small number of model output queries, the framework achieves competitive attack performance with low computational overhead, making it both practical and efficient for real-world scenarios. Extensive experiments on the 3DPW and UBody datasets demonstrate the superiority of LatentStealth, revealing critical vulnerabilities in current systems. These findings highlight the urgent need to address and mitigate security risks in digital human generation technologies.
I Introduction
@@
Expressive human pose and shape estimation (EHPS) from monocular images or videos is fundamental to digital human modeling and key applications such as live streaming, virtual reality, and interactive gaming [24, 28, 43, 39, 2, 12, 11]. In these scenarios, EHPS systems play a critical role in animating realistic digital avatars capable of dynamically replicating human movements and facial expressions in real time, as shown in Fig. 1 (a). Although recent advancements have significantly enhanced estimation accuracy, the security vulnerabilities of such models remain largely underexplored. Growing evidence [36, 3, 31, 42] indicates that these systems are susceptible to adversarial perturbations. Although current attack methods have demonstrated promising effectiveness, their perturbations are typically visible and therefore easily detected, as shown in Fig. 1 (b), which has limited attention to this issue within the research community. However, if attacks are conducted in an imperceptible manner, they may undermine public safety. For instance, in digital human live streaming, such attacks could cause an avatar’s head to appear severed or clothing to fall off, leading to severe incidents involving horror or pornography. Therefore, raising awareness of this problem within both the research community and broader society is urgent.
Figure 1: Motivation and comparison of adversarial attacks on digital human generation. (a) A standard pipeline generates correct postures by disentangling pose, shape, and facial factors. (b) Existing attacks perturb the image space, causing visible artifacts and risky postures. (c) Our method manipulates the latent space, remaining visually imperceptible while still inducing erroneous posture estimation, revealing a more stealthy and realistic attack.
@@
While these models exhibit strong performance, their robustness against adversarial attacks has received limited attention. From a methodological standpoint, existing EHPS studies primarily focus on reducing errors in body pose estimation [43, 39] and facial pose estimation [17, 26, 23, 2]. This is typically achieved using parametric human body models such as SMPL-X [28] or SMPLer-X [2], which leverage deep neural networks to capture the complex geometry and motion of the human body. TBA [22] was the first to propose a white-box transferable adversarial attack against EHPS. It utilizes the gradient of the SMPLer-X model to generate adversarial examples that significantly increase the estimation error across various EHPS models, thereby exposing potential security vulnerabilities. However, the perturbations produced by TBA are visually perceptible and fail to pass human inspection, which limits its applicability in real-world scenarios. Despite its effectiveness, such attacks have yet to attract sufficient attention from the research community. Considering these factors, we argue that an ideal adversarial attack on EHPS should meet three key criteria: i) the perturbation must be visually imperceptible; ii) the computational cost of the attacker should be minimal, aligning with realistic threat models; iii) the adversarial examples must achieve a high error growth rate. A substantial gap remains between current methods and these ideal standards.
@@
To tackle this challenge, we propose a novel unnoticeable attack method, termed LatentStealth, against EHPS models. Inspired by [6, 30], LatentStealth perturbs the latent space of a pretrained Variational Autoencoder (VAE) [19, 10] by injecting low-magnitude Gaussian noise. Leveraging the nonlinear generative capabilities of the decoder of VAE [18, 34], our proposal can produce an initial adversarial example that is visually indistinguishable from the original input, as shown in Fig. 1 (c). The resulting pixel-wise difference is treated as an initial perturbation and subsequently refined through an iterative optimization process guided by a customized multi-task loss function. This loss simultaneously maximizes the discrepancy in the EHPS model’s predictions while minimizing perceptual deviation from the clean image. Moreover, since the attack operates under a strict test budget, its computational cost remains low, ensuring efficiency. We conduct extensive experiments on the 3DPW and UBody datasets to evaluate the effectiveness of LatentStealth. Results show that it achieves strong attack capability (increasing pose estimation errors by an average of 17.27%–58.21%), high stealthiness (achieving the best FID [9] and PSNR [38]), and low computational cost (requiring only three queries of the model outputs).
In conclusion, our main contributions are:
  • We propose the first unnoticeable adversarial attack method for EHPS systems under realistic constraints, which significantly degrades model performance while remaining visually undetectable. Moreover, our study highlights a critical yet underexplored security risk in digital human modeling.
  • We develop LatentStealth, a VAE-based attack that perturbs latent representations with low-magnitude noise and leverages the decoder’s nonlinear generative capacity to craft adversarial examples indistinguishable from clean inputs. This design fundamentally differs from pixel-space attacks by enhancing both stealthiness and naturalness.
  • We design a multi-task loss that jointly balances adversarial effectiveness and perceptual similarity. Coupled with an iterative refinement process constrained by a strict test budget, LatentStealth achieves strong attack performance with low computational overhead, ensuring practical applicability in real-world scenarios.
The remainder of this paper is organized as follows. Section II reviews the related work. Section III introduces the necessary background knowledge. Section IV details the proposed method. Section V reports comprehensive experimental results and comparisons with existing methods. Section VI concludes the paper. Section VII represents the limitation and future work. Section VIII discusses social impacts.
II Related Work
II-A Expressive Human Pose and Shape Estimation (EHPS)
@@
EHPS is fundamental to digital human modeling, with applications in virtual environments, live broadcasting, avatars, and animation [37, 21, 29, 44, 27, 32]. Recent years have seen significant advances in improving the accuracy and realism of EHPS models. [28] introduced SMPL-X, a comprehensive 3D model incorporating articulated hands and expressive facial features for improved monocular pose capture. [26] proposed Hand4Whole, leveraging metacarpophalangeal joints to predict wrist rotations precisely while excluding full-body context for better finger estimation. [23] developed OSX, a one-stage framework using a component aware transformer to reconstruct whole-body meshes without post-processing. [2] presented SMPLer-X, a generalist foundation model trained on diverse datasets, achieving strong generalization and competitive performance across EHPS benchmarks. However, existing research has predominantly concentrated on enhancing estimation accuracy, with limited attention to the models’ vulnerability under attack threats.
II-B Adversarial Attack against Deep Neural Networks.
@@
Adversarial attacks are commonly employed to assess the robustness of deep neural networks (DNNs) against malicious inputs [15, 8, 32, 13, 14, 41]. These attacks introduce subtle perturbations that mislead models while preserving the visual appearance of the input. [16] proposed TT3D, which integrates NeRF-based multi-view reconstruction with dual optimization of texture features and MLP parameters to generate natural-looking, transferable 3D adversarial examples. [5] presented the first content-aware, unrestricted adversarial attack based on diffusion models, enhancing transferability and realism by optimizing semantic perturbations in latent space. [4] proposed DiffAttack, a diffusion-based framework that achieves highly transferable and imperceptible attacks by jointly modulating cross-attention and self-attention mechanisms. Despite these advancements, the robustness of EHPS models remains largely underexplored. [22] proposed TBA consisting of a dual heterogeneous noise generator, the first EHPS-targeted attack; however, its visible perturbations limit its practicality. Consequently, TBA has failed to draw sufficient attention from the research community regarding EHPS security risks. To address this gap, we propose an unnoticeable attack method that reveals critical vulnerabilities in EHPS models.
@@
Figure 2: Overview of the LatentStealth pipeline. The framework consists of two stages: a) noise injection in latent space, and b) noise enhancement.
III Preliminaries
The Main Pipeline of EHPS. Expressive human pose and shape estimation (EHPS) aims to reconstruct detailed 3D representations of the human body, face, and hands from monocular images. This capability underpins a wide array of applications, including virtual environments, live broadcasting, avatar generation, and animation. Formally, the ground-truth pose is denoted as $\alpha\in\mathbb{R}^{53\times 3}$, comprising the body pose $\alpha_{\text{body}}\in\mathbb{R}^{22\times 3}$, left- and right-hand poses $\alpha_{\text{lhand}},\alpha_{\text{rhand}}\in\mathbb{R}^{15\times 3}$, and the jaw pose $\alpha_{\text{jaw}}\in\mathbb{R}^{1\times 3}$. Additionally, individual-specific geometry and facial expressions are encoded by the shape vector $\beta\in\mathbb{R}^{10}$ and the expression vector $\gamma\in\mathbb{R}^{10}$, respectively. Let $\mathcal{P}$ denote the EHPS model; given a unseen monocular image $x$, the objective is to estimate the complete set of pose, shape, and expression parameters $(\alpha,\beta,\gamma)$, e.g., $\mathcal{P}(x)=(\alpha,\beta,\gamma)$. The estimation process minimizes the parameter reconstruction loss:
$\min_{{\alpha,\beta,\gamma}}\left(\|\alpha-\alpha_{gt}\|_{2}+\|\beta-\beta_{gt}\|_{2}+\|\gamma-\gamma_{gt}\|_{2}\right).$ (1)
Finally, 3D joints are rendered using the SMPL-X function $\mathcal{S}$ and a joint regressor $\mathcal{R}$, denoted as $\mathcal{R}(\mathcal{S}(\alpha,\beta,\gamma))$. $(\alpha_{gt},\beta_{gt},\gamma_{gt})$ denotes the ground-truth human pose and shape parameters.
Threat Model. This paper investigates adversarial attack techniques targeting EHPS models. Specifically, the attacker lacks access to any details of the model’s training process, including the dataset, architecture, and trianing trick, and is only permitted a limited number of acquring the model’s output. The generated adversarial examples are directly submitted to the model for inference. Other attack types with limited practical applicability (e.g., training-controlled or model modification attacks) require additional attacker capabilities and are beyond the scope of this study.
IV Method
IV-A Problem Formulation
As noted in previous sections, existing EHPS models primarily aim to reduce human pose and shape estimation errors, with limited emphasis on their vulnerability to attacks. Although prior work [22] introduced the first adversarial attack against EHPS, exposing its inherent fragility, its real-world applicability remains constrained due to the visible nature of adversarial perturbations. This limitation hinders the community’s full recognition of the intrinsic risks associated with EHPS. To tackle the challenge, we propose an unnoticeable attack method, termed LatentStealth, designed for realistic scenarios where the attacker can only access the model’s output under a strict query budget.
Attacker’s Goals. The LatentStealth is designed with two primary objectives: 1) effectiveness and 2) unnoticeability. Specifically, the effectiveness demands that the adversarial examples significantly increase errors in human pose and shape estimation, and the computational cost remains low. The unnoticeability ensures that EHPS model owners are unable to identify the adversarial examples. The formalization is as follows:
$\mathcal{P}(\hat{x})=(\hat{\alpha},\hat{\beta},\hat{\gamma}),\quad\mathcal{P}(x)=(\alpha,\beta,\gamma),$ (2)
$\begin{split}\max_{\hat{\alpha},\hat{\beta},\hat{\gamma}}\;&\Bigl(\|\hat{\alpha}-\alpha\|_{2}^{2}+\|\hat{\beta}-\beta\|_{2}^{2}+\|\hat{\gamma}-\gamma\|_{2}^{2}\Bigr)\\ \text{s.t.}\;&\|\hat{x}-x\|_{2}^{2}\leq\varepsilon,\end{split}$ (3)
where $x$ denotes a unseen monocular image (clean input image), $\hat{x}$ represents the corresponding adversarial sample image, and $\varepsilon$ indicates the perturbation threshold.
IV-B LatentStealth
Existing EHPS attacks often produce visible perturbations, limiting practicality. To overcome this, we propose LatentStealth, an unnoticeable attack method (as shown in Fig. 2) that perturbs the latent space of a pretrained Variational Autoencoder (VAE) [19] with low-magnitude noise and refines it via a multi-task loss under a strict test budget. LatentStealth achieves effective, efficient, and visually stealthy attacks, exposing critical yet overlooked security vulnerabilities in EHPS systems.
@@ IV-B LatentStealth
Noise Injection in Latent Space. Inspired by [33], who questioned the necessity of norm-bounded pixel constraints, we introduce adversarial perturbations in the latent space, thereby more naturally preserving the structural features of the input and improving the imperceptibility of the perturbations. Unlike conventional methods that manipulate input images directly [7, 25, 5, 4], this strategy targets internal representations that more directly govern model predictions. Importantly, perturbing the latent space allows for adversarial behavior while preserving high perceptual fidelity in the image space. To reduce attack cost and improve real-world applicability, we employ a publicly available pretrained VAE [19] without additional training or fine-tuning.
@@
The VAE is trained by maximizing the evidence lower bound and utilizes the reparameterization trick to enable differentiable sampling as follows:
[…]
@@
where $\phi$ and $\theta$ are the encoder and decoder parameters, the encoder outputs $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$ as mean and standard deviation, $z\sim\mathcal{N}(0,I)$ is the standard Gaussian prior, and the decoder $\mathcal{G}_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ maps the $d$-dimensional latent space to the $n$-dimensional pixel space, yielding a smooth latent representation [20].
[…]
@@
[…]
@@
$\begin{split}\mathcal{P}(\hat{x})&=\mathcal{P}\!\big(x+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\big)\\ &\approx\mathcal{P}(x)+J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).\end{split}$ (10)
We define the threshold for significant changes in the estimation parameters of the EHPS model as $\delta$. If the deviation exceeds this threshold, the model is considered to be misled, i.e., $\mathcal{P}(\hat{x})-\mathcal{P}(x)||_{2}\geq\delta$. According to Eq. 10, this leads to the inequality $J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\geq\delta$.
By combining the Jacobian matrices into a single term $\mathbb{J}=J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)$, the inequality can be reformulated as $\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||_{2}\geq\delta$. Given that $\epsilon_{2}\sim\mathcal{N}(0,I)$, the expected perturbation is:
[…]
@@
$\mathbb{J}\cdot(\eta\cdot\upsilon_{1})||_{2}=\eta\cdot\sigma_{1}\geq\delta\Rightarrow\eta\geq\frac{\delta}{\sigma_{1}}.$ (13)
@@
$\begin{split}\mathcal{P}(\mathcal{G}_{\theta}(z+\Delta{\hat{z}}^{(t)}))&=(\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),\\ \mathcal{P}(x)&=(\alpha,\beta,\gamma),\end{split}$ (14)
[…]
@@
$\nabla_{\Delta{\hat{z}}^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{G}_{\theta}}\frac{\partial\mathcal{G}_{\theta}}{\partial\Delta{\hat{z}}^{(t)}},$ (16)